OpenStudy (goldrush18):

The function f on R is defined by F(X)={x-3 if x<=3 {x/4 if x>3 Find the value of i. f [f(20)] ii. f [f(8)] iii. f [f(3)]

OpenStudy (experimentx):

okay ... i'll do the first example ... you do the rest 20 > 3 then F(x) = x/4 .... right?? f(20) = 20/4 = 5 again, 5>3, so F(x) = x/4 ... so f(5) = 5/4 = 1.25 :D

OpenStudy (experimentx):

try no 2, and alert me if you are stuck ..

OpenStudy (goldrush18):

ok im gonna try it

OpenStudy (goldrush18):

f(8)=x/4 =8/4 =2 so 2<3 so f(x)= 2-3= -1?

OpenStudy (experimentx):

yes ... you are right!!

OpenStudy (goldrush18):

omg i got it :)

OpenStudy (goldrush18):

thanks

OpenStudy (experimentx):

yw